# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v4

@@ 4.1 Experimental Setting
Evaluation Metrics. We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”. “Transfer”evaluates the model’s zero-shot transfer capability on unseen data. “Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last” evaluations.
Implementation Details. We follow the experimental setups in [63, 58, 55] and use the CLIP model with a ViT-B/16 backbone [40] for all experiments. By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [31]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold adaptive pruning is fixed at 0.005 for both settings.
@@ Methods
Table 3: Comparison of computational cost with prior methods.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the X-TAIL setting evaluation results in Table 1. In contrast to prior methods [58, 55] [^3] The DUAL version of [55] stores all the training samples’ feature embeddings in memory. For a fair comparison, we exclude these results here and provide more detailed comparisons in the Appendix., which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
[…]
4.3 Discussions
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [40] 68.24 33.77 66.72 56.24
MoE-Adapter [58] 68.24 33.77 66.72 56.24
RAIL [55] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Table 4: Evaluation of average transfer accuracy on unseen datasets not encountered during continual learning.
Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset using our proposed method.
Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset using our proposed method.
@@ 4.3 Discussions
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and domain adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix B.3.
@@
Figure 7: Ablation studies on different insertion locations and the impact of varying the initial rank.
Figure 8: Analyses of the effects of (a) varying the maximum pruning threshold and (b) different ratios of dense training iterations.
@@ 4.4 Ablation Studies
Effect of maximum threshold value and dense iteration ratio. We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio of dense training iterations in Fig. 8. Our results indicate that as the threshold increases, Last accuracy improves, suggesting that pruning more ranks can enhance transfer performance on unseen data. However, aggressive pruning limits adaption ability to to downstream tasks, as reflected in changes to the Last accuracy metric. We find that setting the threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
The dense training iteration ratio determines when adaptive pruning begins. Setting this ratio to 0 (i.e., pruning from the start of training) hinders adaptation, as the model lacks sufficient time to learn task-relevant knowledge before pruning begins. Conversely, delaying pruning too long increases the risk of forgetting pre-trained capabilities, ultimately reducing transfer accuracy. We find that setting this ratio between 0.5 and 0.7 achieves the best trade-off across all experimental settings, effectively balancing both adaptation and knowledge retention.
@@ 5 Conclusion
Limitations and future works. Our approach merges updates into pre-trained weights without modeling inter-task correlations. Future works could explore leveraging inter-task relationships and extending our method to other VLMs.
Acknowledgment
@@ Appendix A Additional Details of the Proposed Method
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
@@ B.4 Evaluation with Other VLMs
Table 6: Averaged performance using BLIP.
@@ B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
Table 7: Performance of 16-shot PEFT on downstream tasks.

### v5

@@ 4.1 Experimental Setting
Evaluation Metrics. We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”. “Transfer”evaluates the model’s zero-shot transfer capability on unseen data. “Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last”.
Implementation Details. We follow the experimental setups in [66, 61, 58] and use the CLIP model with a ViT-B/16 backbone [42] for all experiments. By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [33]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold for adaptive pruning $\kappa$ is fixed at 0.005 for both settings.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the X-TAIL setting evaluation results in Table 1. In contrast to prior methods [61, 58] , which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
[…]
4.3 Discussions
Table 3: Average transfer accuracy on unseen datasets not encountered during CL.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [42] 68.24 33.77 66.72 56.24
MoE-Adapter [61] 68.24 33.77 66.72 56.24
RAIL [58] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
Table 4: Computation cost.
@@ Methods
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix.
@@
Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
Figure 8: Analyses (a) maximum pruning threshold and (b) ratios of dense training iterations.
@@ 4.4 Ablation Studies
Analyses of pruning threshold and dense-iteration ratio. To better understand the trade-off between downstream learning and forgetting, we control the number of adaptive ranks by varying the maximum threshold $\kappa_{\max}$ and the fraction of dense training iterations (Fig. 8). Gradually raising $\kappa_{\max}$ improves Last accuracy by pruning more useless or noisy ranks, whereas overly aggressive pruning limits task adaptation; a modest setting of 0.005 strikes a good balance. The dense-iteration ratio governs pruning onset: pruning from the start (ratio = 0) removes ranks before it can learn from the training data, while delaying pruning too long increases the risk of retaining useless or noisy ranks that do not get pruned eventually.
@@ 5 Conclusion
Limitations and future works. Our approach merges updates into pre-trained weights without modeling inter-task correlations, mainly evaluated using pre-trained CLIP model. Future works could explore leveraging inter-task relationships and extending our method to other PTMs and settings.
Acknowledgments and Disclosure of Funding
@@ Appendix B More Experimental Results and Ablations
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper. We report experiment results averaged over 3 independent runs for reproducibility. Experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
@@ B.4 Evaluation with Other VLMs
Table 7: Performance of 16-shot PEFT on downstream tasks.
@@ B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
Broader Impacts
CoDyRA’s sparse, adaptive, importance-aware low-rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on-device personalization in domains such as mobile health and education and empowering resource-constrained users and organizations. As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors. We therefore recommend transparent update logging, regular audits, and differential-privacy safeguards. Although more efficient than full fine-tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.